PURE ROTATIONAL AND FINE-STRUCTURE TRANSITIONS OF NITRIC OXIDE (NO) IN THE FAR-INFRARED

\underline{THOMAS D. VARBERG}, Department of Chemistry, Macalester College, St. Paul, MN 55105; FRED STROH, Institute for Stratospheric Chemistry, Forschungszentrum Jülich GmbH, D-52425 Jülich, Germany; KENNETH M. EVENSON, Time and Frequency Division, National Institute of Standards and Technology, Boulder, CO 80303.

The absorption spectrum of the ^{14}NO and ^{15}NO isotopomers of nitric oxide in the far-infrared has been measured and analyzed. Pure rotational transitions up to J′′ = 37-1/2 within the ^{2}Π_{1/2} and ^{2}Π_{3/2} spin components and several 2Π_{3/2} ← ^{2}Π_{1/2} fine-structure transitions were recorded within the ground vibrational state. A least-squares fit to these data combined with lambda-doubling transitions measured by previous workers has resulted in significant improvement in the accuracy of the rotational and fine-structure parameters of these two isotopomers. Improved values for the lambdadoubling and hyperfine constants will also be reported.